Calculus I, Fall 2013 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 04

1. Section 3.6: pick any 5 from problems 51, 53, ⋯, 75.

2. Assume g(2) = 3, g′(2) = 0.1, f′(2) = 3, f′(3) = 4 and f′(4) = 5. What is \frac{d}{dx}f(g(x)) at x = 2?

3. Section 3.7: problems 27, 31, 42.